OpenStudy (anonymous):

5/7÷2/7=?????? A 6/7 B 2 1/2 C 10/49. D 2/5

OpenStudy (anonymous):

\[\frac{ 5 }{ 7 }\div \frac{ 2 }{ 7 }=\frac{ 5 }{ 7 }\times \frac{ 7 }{ 2 }=?\]

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

Help

OpenStudy (blurbendy):

5/7 divided by 2/7 is the same as 5/7 times 7/2 It's a rule that can help you save a lot of time

OpenStudy (anonymous):

35/14

OpenStudy (anonymous):

Thanks guys

OpenStudy (blurbendy):

np